# Foundation Models for Structural Health Monitoring

### v1

@@ 4 Structural Health Monitoring Installations and Datasets
This section describes the three SHM use cases considered in our work.
4.1 Use Case 1 (UC1): Anomaly Detection Task
The first use case exploits the vibration data collected on a dual-carriageway state highway located in northern Italy. The two carriageways are on two independent viaducts, composed of 9 different sections each. In 2019, one of the spans of the viaducts underwent a technical intervention to strengthen the structure, which caused a change in the vibration signal collected. Before the intervention, this section was instrumented with five SHM nodes, directly placed under the bridge to monitor viaduct vibration. Therefore, vibration data from both before and after the intervention were gathered from the sensors. These data are used in this study as a proxy for an abrupt change in the viaduct structure caused, for example, by external factors such as an earthquake. Namely, we consider vibrations raw data collected after the intervention as the normal data produced by a “sane” viaduct, while the accelerations measured before the intervention as the “anomaly”, given the high degradation of the bridge’s structure. Sensor nodes for vibration monitoring in this installation use the LIS344ALH analog 3-axial accelerometer [42] with a full-scale range of ±$\pm 2/\pm 6g$, that gather vibration data in the three directions, x, y, z, with an angle of $90^{\circ}\pm 2\%$ between each other. The system also includes an HTS221 temperature and humidity sensor [43], and an STM32L476VGTx microcontroller as a computational core, used for data analytics tasks and the scheduling of the data collection steps. The microcontroller unit is an ARM 32-bit Cortex-M4 running at 80 MHz, with 96 KB of SRAM and 1MB of Flash memory.
Dataset
The employed dataset is the same described in [36], comprising a total of 15 days of continuous monitoring of the viaduct. For our analysis, we consider only the central sensor of the chain, which is the one most influenced by the viaduct vibration. The data include measurements from 4 days before the maintenance intervention, labeled as anomalies, and two intervals of 4 and 7 days respectively, after the intervention, labeled as normal data. We select as the test set the two intervals of 4 days to have a perfect balance between anomalies and normal data. The remaining 7 days of normal data are used to train the AD algorithm. Note that anomalies are not used during training given that, for this task, we only consider one-class models trained to identify anomalies as outliers. The latter are, therefore, used only at test time to assess the detection accuracy.
4.2 Traffic Load Estimation Task
4.2.1 Use Case 2 (UC2)
As a first benchmark for TLE, we consider another viaduct located in Italy, equipped with a SHM sensor network since 2017. A view of the viaduct and of the installed sensors is given in Fig. 2. The viaduct is sustained by five concrete pillars and has a total length of $\sim$580 m. The spans are 67 m-112 m long. Each span is formed by a steel caisson inside which there are 12 prestressing cables. The height of the cross-section of the main beam varies from 6.0 m (at the bearings) to 3.0 m (at the center line of each span). After the construction, a series of unbound tendons (27 strands each) were placed, prestressed, and anchored to the abutments. Since the structural safety of a prestressed bridge is highly dependent on the durability of these cables, a system of sensors (see Fig. 2) has been installed on them to check for breakages. Since September 2017, the sensor network has been fully active and is gathering data stored in the cloud. The network comprises 90 sensors that gather data from all the spans positioned in the upper part of the pre-stressed cables. Each node comprises an STM32F405RG microcontroller (MCU), which takes care of managing the data acquisition and optionally performs simple data analytics steps, and the same accelerometer, temperature and humidity sensors described in Section 4.1.
Dataset
We consider the data of a single section, specifically section 10, which corresponds to the area of the road in which an optical camera has been installed for the labelling process, i.e., to compute the ground truth reference for TLE. This section comprises 7 sensors, mounted on different tendons. Our study uses only the central one, which experiences the highest vibrations. In detail, we consider 31 minutes of recorded accelerator from the z-axis of the sensor and a synced video from the camera. Train and test data are obtained considering a 70:30 split ratio: training data consists of 21 minutes and 42 seconds, while the other 9 minutes and 18 seconds are used for testing. Data were collected between 8.00 a.m. and 9.00 a.m., with the camera working at 10 Frames Per Second (FPS). Labels were extracted from the video assigning value 1 to acceleration samples corresponding to each camera frame in which a light vehicle crossed an ideal vertical line placed above the pillar of section 10, and value 2 for each crossing by a heavy vehicle. Value 0 was associated to frames without crossing vehicles. Noteworthy, given the sampling frequency of the camera (10 FPS) and accelerometer (100 Hz), slices of 10 consecutive acceleration samples always receive the same label. Acceleration data and labels were then synchronized to form a single multi-variate time series dataset, with $100\times 1860s$ (i.e., 31 minutes $\times$ 60) samples. The final TLE label was computed as the sum of the 1s and 2s in the window divided by 10, as in [10] (details in Sec. 5.1).
4.2.2 Use Case 3 (UC3)
The third case study involves a roadway bridge during normal operation. The bridge consists of 18 spans and is 583 meters long, with two lanes. It is a reinforced concrete girder bridge with an isostatic static scheme. Each span is 20 meters long, except for the first and last, which are 10 and 29.5 meters long, respectively. An industrial SHM system has been recording data since the beginning of 2022. It includes 282 MEMS biaxial inclinometers and 142 MEMS triaxial accelerometers with a $\pm$2,g full scale and 100 Hz sample rate. They are evenly distributed across spans, with accelerometers monitoring the two external beams in each span. Three sensors are installed for each beam at the quarter, third, and midspan positions.
In addition to the monitoring system, a WiM system is located about 600 meters before the viaduct. This section does not have any highway exits or parking areas. The WiM system provides a different source for ground truth data for the study, compared to Use Case 2. Namely, it offers detailed information about the traffic on the bridge, including the lane of the detected vehicle, its length, weight, speed, and the number of vehicle axles. The ground truth labels from the WiM are associated with data gathered by the sensors by applying a time offset equal to the distance between each sensor and the WiM system, divided by the vehicle’s speed.
Dataset
We utilize two days of data to train and evaluate our algorithms. We use the first complete day’s data, spanning from 00.00 to 23.59, and covering a variety of traffic conditions, only as pre-training data for the foundation model. We then use the first half of the following day (00.00 to 12.00) as additional training data for supervised fine-tuning on this specific task. Finally, the last half day’s data is reserved for testing. It’s important to note that, in the 600 m between the WiM and the bridge, vehicle speeds may vary, with potential lane changes, vehicle overtaking, and overlapping. As a result, vibrations on the bridge may not be entirely correlated with the data collected from the WiM, which can unavoidably lead to a loss of accuracy in this analysis.
5 Methodology
This section presents our novel approach to vibration-based SHM leveraging self-supervised learning and Transformer foundation models. Firstly, we outline the data processing toolchain and train/test splitting (Section 5.1). Then, we delve into our model architecture (Section 5.2) and training strategies (Section 5.3). The core objective of our work is to understand whether it is possible to pre-train a large Transformer model on substantial amounts of unlabeled data and then fine-tune it to realize different SHM tasks. We consider both AD and TLE as downstream tasks to validate this hypothesis, focusing on all three Use Cases (UCs) described in Sec. 4. Specifically:
5.1 Pre-processing Pipeline
The complete data processing pipeline used in our work is shown in Figure 3. This section outlines the pre-processing applied before inputting data into our masked autoencoder. The pre-processing settings are consistent for training and validation data across all analyzed UCs, except when expressly indicated.
The first pre-processing step is windowing, i.e., the division of the time-series data into time windows, which the algorithm can process. For UC1, we use 5s windows, as in the reference paper [36], with a stride of 2s to increase the number of training samples. For the other use cases, we enlarge the window to 60s since this dimension yields superior performance in the reference study [10], using a 2s stride in UC2, and 15s stride, given the very large data size, in UC3. All windows are composed of uni-dimensional signals since we always employ either a single sensor or multiple sensors, but consider them separately. We also only consider the z-axis of the acceleration data. Thus,each window is always a single vector of dimension $T=100\times T_{win}$, where 100 is the acceleration sampling frequency in all three datasets. All time windows are normalized with mean and standard deviation.
After creating the windows, we apply energy filtering, i.e., we disregard as irrelevant those windows whose cumulative energy is under a certain threshold, $th=3.125\times(10^{-5})$ for UC1 and $th=1.25\times 10^{-6}$ for UC2 and UC3. Energy thresholds are computed using the iterative process proposed in [36]. Then, since masked autoencoders have been shown to reconstruct signals better in the frequency domain [17], the acceleration spectrogram is extracted. Precisely, we feed the model with 100x100 time-frequency representations in the form of spectrograms. The first dimension corresponds to the temporal domain, while the second captures the frequency domain. While we notice that this window is big enough to obtain good task performance, we did not explore the downscaling of the spectrogram image to reduce the models’ computational complexity, as the reduction of the input size has an impact only on the dimension of the last layer, whose size is negligible with respect to the whole model size.
$y=\frac{\sum_{t=1}^{T}(l_{t}=k)}{10}$ (1)
5.2 Foundation Model Architecture
Figure 4: Architecture of our masked autoencoder during pre-training.
Our model architecture draws inspiration from the Masked Autoencoder proposed in [17], which was originally introduced to learn self-supervised representations from audio spectrograms. The model uses a stack of standard Transformer blocks for both the encoder and the decoder, each consisting of multiple attention layers. As a backbone, we use ViT-Base (ViT-B), but we simplify it by employing only three attention blocks for the encoder and decoder, drawing inspiration from the smallest option in [44]. The model takes as input acceleration spectrograms divided into non-overlapped regular grid patches, with a configurable portion of them being masked during training. The complete architecture is shown in Fig.  4.
To fine-tune our foundation model on the considered downstream tasks, architectural modifications are necessary. Namely, the model architecture remains unchanged when fine-tuning on UC1, because the task is framed as an outlier detection [36]: during training, the model learns the latent representation of normal data only, so that during testing, the decoder is expected to fail in accurately reconstructing data that deviates from its learned patterns. Accordingly, those samples whose reconstruction error (computed as detailed in Sec. 5.3.1) exceeds a user-defined threshold are classified as anomalies or outliers. On the other hand, the decoder is removed when fine-tuning on TLE tasks, and a single fully connected layer is appended directly to the latent representation produced by the encoder. The added layer has a single output neuron and is trained to predict the scalar target traffic variable defined in Eq. 1.
5.3 Training Phases
5.3.1 Pre-training
The full encoder-decoder model, with the architecture of Fig. 4 is initially pre-trained on a large amount of unlabeled data. We select pre-training data following the split described Section 2, combining the training sets of all three UCs. Note that, although our data have labels, we do not use them in this phase, to simulate a scenario where the costly process of labelling all available samples is not affordable.
During pre-training, the non-overlapping spectrogram patches of each input sample are masked with probability $p=0.8$ (masking ratio) and aligned, before being passed to the transformer encoder. The latter then processes the non-masked patches to produce latent representations. Encoded patches are padded with trainable mask tokens and fed to the decoder, whose objective is to reconstruct the original spectrogram, including the missing parts. The encoder and decoder are jointly optimized to minimize the Mean Squared Error (MSE) between the reconstruction and the input spectrogram, averaged over masked patches. This pre-training is completely self-supervised and task-independent. Its objective is to learn robust latent representations of SHM data that allow the model to infer the underlying patterns in the data despite the missing information.
5.3.2 Fine-tuning
The details of the fine-tuning phase differ depending on the downstream task. For UC1 (AD), as mentioned above, we maintain the autoencoder structure of our model and use a threshold on the reconstruction error to identify anomalies. Therefore, since the fine-tuning architecture is the same as pre-training, the model is further trained in an unsupervised way using the same procedure just described, but considering only normal data from the UC1 dataset.
For UC2, the pre-trained model is fine-tuned considering the 21 minutes of coupled acceleration data and video recording on the viaduct presented in Sec. 4.2.1. Two separate fine-tunings are executed for light and heavy vehicle TLEs, starting from the same pre-trained model. For UC3, we consider the first half of the second day’s data to fine-tune our foundation model, as described in 4.2.2. For these two downstream tasks, we follow a classic supervised fine-tuning procedure, where the model learns to predict the traffic load values rather than to reconstruct the input spectrograms. As a loss function, we use the MSE between the newly inserted fully connected layer’s output prediction and the ground truth target variable.
5.4 Model Size Exploration and Knowledge Distillation
We carry out an extensive search on the optimal size of our foundation models. To this end, we systematically decrease the embedding dimensions of both the encoder and the decoder from the baseline dimensions $(e_{dim},d_{dim})=(768,512)$ scaling them progressively by a factor $2$ down to $(e_{dim},d_{dim})=(24,16)$. These new models are pre-trained on all datasets and subsequently fine-tuned on the three downstream tasks, following the same procedure described in the previous section. Additionally, we considered a modified fine-tuning procedure enhanced with Knowledge Distillation (KD). In this setup, considering the baseline model as a teacher, the student utilizes a multi-objective fine-tuning loss function composed of two contributions: the first one, $\mathcal{L}_{task}$, is the Mean Absolute Error (MAE) computed on student prediction against the ground truth, while the second one, $\mathcal{L}_{KD}$, is the Root Mean Squared Error (RMSE) between student and teacher predictions. Through an empirical approach, we assessed that evenly balancing these two contributions leads to faster convergence of the overall loss, whose function becomes:
6 Results
@@
In this section, we first introduce our comparison baselines and then detail the results of our foundation models on the three use cases. We also quantify the impact of self-supervised pre-training on the performance of our models. Lastly, we analyze the trade-off between model size and performance and compare regular and KD-enhanced fine-tuning. In all experiments except the last one, we consider our largest masked autoencoder version, i.e., $(e_{dim},d_{dim})=$(768, 512).
@@ 6.1 Comparison Baselines
For UC1, we compare against the top-performing model of [36], a PCA that considers a window dimension of 5 seconds (500 samples at 100 Hz) and a post-processing median filter with varying length to aggregate consecutive predictions.
For UC2 and UC3, we compare with the five supervised algorithms proposed in [10], i.e., the work that first introduced the UC2 dataset. Namely, we consider: i) a K-Nearest Neighbors (k-NN); ii) a Linear Regressor (LR); iii) a Random Forest (RF) trained with the variance reduction criterion [47]; iv) a Support Vector Regressor (SVR); v) a Multi-Layer Perceptron (MLP) with ReLU activations, trained with the Adam optimizer, a learning rate of $10^{-3}$ and a mini-batch size of 200. A summary of the state-of-the-art models and their hyperparameters is shown in Table 2.
@@ 6.2 Anomaly Detection (UC1) Results
We evaluate AD models in terms of specificity, computed as the correctly classified normal windows, sensitivity, i.e., the correctly classified anomalous windows, and accuracy, i.e., the ratio of correctly classified windows overall. For both our models and the baseline PCA, we use the same procedure of [36] to determine the reconstruction error threshold associated with anomalies. Namely, the threshold is initialized to the average MSE on normal (training) data, to which we add the standard deviation over a single day of normal data. The value is then progressively increased by a small step until 100% specificity is achieved on this single validation day.
Figure 5 shows the comparative performance of PCA and our foundation model while varying the number of windows’ outputs fed to the median filter as post-processing. The graph highlights the increasing trend of accuracy, sensitivity, and specificity as the dimension increases for both methods. However, we can observe a significant improvement in all evaluation metrics of our model compared to the PCA, especially for short time windows. For example, when applying the median smoothing to 60 windows, our masked autoencoder achieves an accuracy of 99.92%, while the PCA only reaches 75.76%. To match the accuracy of our model, the PCA needs a post-processing median filter of 240 windows, i.e., 4x larger. The sensitivity shows a similar trend: for a 60-windows smoothing, our model reaches 99.9%, compared to 55.68% of the PCA.
@@ 6.3 Traffic Load Estimation (UC2 and UC3) Results
To compare the performance of TLE models, we consider 5 different metrics: MSE, MAE, R2 score, Mean Squared Percentage Error (MSE%) and Mean Absolute Percentage Error (MAE%), where the last two are obtained diving the MSE and MAE by the average predicted value on the test set, allowing to quantify the error magnitude w.r.t. the true values.
@@ 6.3.1 UC2
Table 3 reports the results obtained on UC2 for light and heavy vehicles. Considering the latter, we outperform the best baseline models from [10] (SVR and RF) in all metrics. However, our foundation model reduces the MSE by 3.2x, the MAE by 1.3x, and improves the R2 of +0.06. We impute this to the increased capacity of our foundation model, which permits to better leverage the non-linear and higher-order interactions within acceleration data that correlate with traffic flow, while current SoA machine learning methods exhibit limitations in capturing these intricate relationships. Additionally, notice how there is no single best SoA model, while our approach convincingly overcomes them all in every considered metric. For the prediction of the light vehicle traffic, our foundation model outperforms the SoA in terms of MSE (1.6x reduction) and R2 (+0.06) and is the second best in terms of MAE (1.03x increase over the best baseline model, the RF). Importantly, having a significantly lower MSE and a slightly higher MAE means that our approach does not output predictions that are far off the ground truth, as the quadratic function of the MSE penalizes major mistakes. Thus, it could be more amenable for a real-world deployment than the baseline RF.
@@ 6.3.2 UC3
The results of UC3 are shown in Table 4. We compare our approach with the same models considered for UC2. However, since there is no previous literature on this dataset, we re-train all baseline models with the same data split used for our foundation model while keeping the same hyper-parameter settings of [10]. Results show that our foundation model overcomes state-of-the-art performance on all metrics. The second-best performing model, the MLP, fails to reach acceptable results, demonstrating the higher complexity of this UC. Our fine-tuned foundation model can decrease the MSE and MAE of the MLP by 1.9$\times$ and 1.5$\times$, respectively, while increasing the R2 score by 5.4$\times$ as well.
It is worth noticing that this is the task with the largest dataset available for training: this could further explain the superiority of our approach, which is expected to thrive when more data are available, and, at the same time, justify the limits of SoA approaches, as their shallow nature may be unsuited to handle this complexity. The observed performance degradation for both state-of-the-art and our proposed approaches on UC3 compared to UC2 can be attributed to two key factors. Firstly, sensor placement may be a factor influencing task difficulty. In fact, unlike UC2, sensors are positioned on two external beams per span, which might lead to reduced sensitivity to vehicle-induced vibrations. Secondly, as anticipated in Section 4, the label assignment procedure likely introduces noise into the training data. The physical separation of the WiM system from the bridge disrupts the correlation between bridge vibrations and WiM collected data. While the model might be accurately predicting traffic load based on the collected data, the ground truth labels assigned based on the WiM system may not accurately reflect real-world bridge behaviour. This misalignment between predicted and actual bridge conditions may partially explain the performance drop observed on this use case.

### v2

@@ 6 Results
In this section, we first introduce the datasets used for training, validation and testing, summarized in Table 2. Notice that, identically to state-of-the-art works, data splitting is tailored for each task, considering the fundamental differences between anomaly detection (UC1) and TLE (UC2 and UC3). Then, we introduce the comparison baselines and detail the results of our foundation models on the three use cases. We also quantify the impact of self-supervised pre-training on the performance of our models. Lastly, we analyze the trade-off between model size and performance and compare regular and KD-enhanced fine-tuning. In all experiments except the last one, we consider our largest masked autoencoder version, i.e., $(e_{dim},d_{dim})=(768,512)$.
6.1 Datasets
UC1
the employed dataset is the same as the one described in [32], comprising a total of 15 days of continuous monitoring of the viaduct. For our analysis, we consider only the central sensor of the chain, which is the one most influenced by the viaduct vibration. The data include measurements from 4 days before the maintenance intervention, labeled as anomalies, and two intervals of 4 and 7 days respectively, after the intervention, labeled as normal data. We select as the test set the two intervals of 4 days to have a perfect balance between anomalies and normal data. The remaining 7 days of normal data are used to train the AD algorithm (and the AD threshold). Note that anomalies are not used during training given that, for this task, we only consider unsupervised models trained to identify anomalies as outliers.
UC2
we analyze only section 10 of the viaduct, where an optical camera provides ground truth labels for TLE. This section includes seven sensors mounted on different tendons, but we use only the central one, which experiences the highest vibrations. The dataset consists of 31 minutes of recorded z-axis acceleration data and a synchronized video, split 70:30 for training (21 minutes and 42 seconds) and testing (9 minutes and 18 seconds). Data were collected between 8:00 a.m. and 9:00 a.m., with the camera operating at 10 FPS. Labels were assigned based on vehicle crossings, using the video: a value of 1 for light vehicles, 2 for heavy vehicles, and 0 otherwise. Given the accelerometer’s 100 Hz sampling rate and the camera’s 10 FPS, each label was assigned to a set of 10 consecutive acceleration samples. The final TLE label was computed as the sum of 1s and 2s in the window, divided by 10, following [8] (see Sec. 5.1).
UC3
we utilize two days of data to train and evaluate our algorithms. We use the first complete day’s data, spanning from 00.00 to 23.59, and covering a variety of traffic conditions only as pre-training data for the foundation model. We then use the first half of the following day (00.00 to 12.00) as additional training data for supervised fine-tuning on this specific task. Finally, the last half day’s data is reserved for testing. It’s important to note that, in the 600 m between the WiM and the bridge, vehicle speeds may vary, with potential lane changes, vehicle overtaking, and overlapping. As a result, vibrations on the bridge may not be entirely correlated with the data collected from the WiM, which can unavoidably lead to a loss of accuracy in this analysis.
@@ 6.2 Comparison Baselines
We use different baseline algorithms depending on the UC, given the difference in the solved task, i.e., AD vs. TLE. For UC1, we compare against the top-performing model of [32], a PCA that considers a window dimension of 5 seconds (500 samples at 100 Hz) and a post-processing median filter with varying length to aggregate consecutive predictions. We underline that this solution has already been shown to outperform convolutional and fully connected deep autoencoders, as demonstrated in [32].
For UC2 and UC3, we compare with the five supervised algorithms proposed in [8], i.e., the work that first introduced the UC2 dataset. It is worth noticing that this is the task with the largest dataset available for training: this could further explain the superiority of foundation-model-based approaches, which are expected to thrive when more data are available, and, at the same time, justify the limits of SoA approaches, as their shallow nature may be unsuited to handle this complexity. Namely, we consider: i) a K-Nearest Neighbors (k-NN); ii) a Linear Regressor (LR); iii) a Random Forest (RF) trained with the variance reduction criterion [43]; iv) a Support Vector Regressor (SVR); v) a Multi-Layer Perceptron (MLP) with ReLU activations, trained with the Adam optimizer, a learning rate of $10^{-3}$ and a mini-batch size of 200. In addition to that, we consider two alternative deep learning architectures, a Temporal Convolutional Network (TCN) and a Long Short-Term Memory (LSTM) network. We implement the TCN architecture of [44] composed of four 1D convolutional layers with increasing dilation factor from first to last block, followed by a regression head consisting of two linear layers with ReLU activation and dropout. For the LSTM, we consider a variation of the architecture from [45] reducing its layers from 4 to 3 due to overfitting, and changing the hidden size to match the one of our Transformer Encoder (768). Also in this case, the regression head is composed of two linear layers with ReLU activation and dropout. On these last two architectures, we test both direct supervised training on the task-specific dataset, as described in Section 5.3, and a two-phase foundation model approach including self-supervised pre-training on all the un-labeled data followed by supervised fine-tuning. For the latter, we use the same masked autoencoder framework of Figure 4, except that the Transformer-based Encoder and Decoder are replaced by TCN and LSTM respectively. We build the Decoders architectures symmetrically to the Encoders: for the TCN, we replace convolutions with up-sampling convolutions, while for the LSTM, we add a linear layer to project the output of the encoder to the decoder embedding dimension (512), once again matched to the one of our Transformer. Section 6.4 only reports the results with pre-training, which has proven superior for both types of model. A summary of the baseline models and their hyperparameters is shown in Table 3.
@@ 6.3 Anomaly Detection (UC1) Results
We evaluate AD models in terms of specificity, computed as the proportion of correctly classified normal windows, sensitivity, i.e., the proportion of correctly classified anomalous windows, and accuracy, i.e., the overall ratio of correctly classified windows, providing a general measure of performance. Figure 5 shows the comparative performance of PCA and our foundation model while varying the number of windows outputs fed to the median filter as post-processing. The graph highlights the increasing trend of accuracy, sensitivity, and specificity as the dimension increases for both methods. However, we can observe a significant improvement in all evaluation metrics of our model compared to the PCA, especially for short time windows. For example, when applying the median smoothing to 60 windows, our masked autoencoder achieves an accuracy of 99.92%, while the PCA only reaches 75.76%. To match the accuracy of our model, the PCA needs a post-processing median filter of 240 windows, i.e., 4x larger. The sensitivity shows a similar trend: for a 60-windows smoothing, our model reaches 99.9%, compared to 55.68% of the PCA.
@@ 6.4 Traffic Load Estimation (UC2 and UC3) Results
To compare the performance of TLE models, we consider five different metrics: Mean Squared Error (MSE), which measures the average squared difference between predicted and true values; Mean Absolute Error (MAE), which computes the average absolute difference between predictions and ground truth; R2 score, which quantifies how well the model explains the variance in the data, with values closer to 1 indicating better performance; Mean Squared Percentage Error (MSE%), and Mean Absolute Percentage Error (MAE%), obtained respectively by dividing MSE and MAE by the average predicted value on the test set, allowing for a relative error assessment with respect to the true values.
@@ 6.4.1 UC2
Table 4 reports the results in terms of these metrics obtained on UC2 for light and heavy vehicles. Considering the latter, we outperform the best baseline models from [8] (SVR and RF) in all metrics. However, our foundation model reduces the MSE by 3.2x, the MAE by 1.3x, and improves the R2 of +0.06. We impute this to the increased capacity of our foundation model in combination with the transformer architecture, which permits to better leverage the non-linear and higher-order interactions within acceleration data that correlate with traffic flow, while current SoA machine learning methods exhibit limitations in capturing these intricate relationships. Noteworthy, the other deep learning architectures tested do not attain satisfactory performance, being inferior to classical machine learning algorithms Additionally, notice how there is no single best SoA model, while our approach convincingly overcomes them all in every considered metric. For the prediction of the light vehicle traffic, our foundation model outperforms the SoA in terms of MSE (1.6x reduction) and R2 (+0.06) and is the second best in terms of MAE (1.03x increase over the best baseline model, the RF). Importantly, having a significantly lower MSE and a slightly higher MAE means that our approach does not output predictions that are far off the ground truth, as the quadratic function of the MSE penalizes major mistakes. Thus, it could be more amenable for a real-world deployment than the baseline RF.
@@ 6.4.2 UC3
The results of UC3 are shown in Table 5. We compare our approach with the same models considered for UC2. However, since there is no previous literature on this dataset, we re-train all baseline models with the same data split used for our foundation model while keeping the same hyperparameter settings of [8]. Results show that our foundation model overcomes state-of-the-art performance on all metrics, improving the MSE and MAE of the second-best, the LSTM, by 0.03 and 0.02, respectively, while increasing the R2 score by 0.03 as well. We observe two general trends: first, the performance of UC3 is worse compared to UC2’s on every metric and for every algorithm. Second, the second-best performing model is the LSTM, trained with the same foundation-model approach as our transformer. This is in strong contrast with UC2, where both the other deep learning based models, TCN and LSTM, reached low performance.
The observed performance discrepancy for both state-of-the-art and our proposed approaches on UC3 compared to UC2 (and their performance ranking) can be attributed to two key factors. First, as anticipated in Section 4, the label assignment procedure likely introduces noise into the training data. The physical separation of the WiM system from the bridge affects the correlation between bridge vibrations and WiM collected data. This misalignment between predicted and actual bridge conditions may partially explain the performance drop observed in this use case. Sensor placement may also affect task difficulty. In fact, unlike UC2, sensors are positioned on two external beams per span, which might lead to reduced sensitivity to vehicle-induced vibrations. Secondly, UC3 comprises roughly 750,000 samples, over three orders of magnitude more than UC2’s dataset, 1,000 samples. DNN models benefit more than classic ML ones from large-scale data, reaching, as in the case of the LSTM, significantly better performance compared to UC2.